Three travelers arrived at a town where there are five hotels. In how many ways can they select their room each at a different hotels?
Answer by ikleyn(52778)

.
5*4*3 = 60 ways.

The first traveler can be accommodated in 5 ways (in any of 5 hotels)

The second traveler can be accommodated in 4 ways (in any of remaining hotels.

The third traveler can be accommodated in 3 ways (in any of 3 = 5 - 2 hotels).

In the rest, these selection are independent.

Therefore, there are 5*4*3 = 60 ways to accommodate, in total.